A ball is shot out of a cannon with its flight modeled by f(x)=-16t2+80t. t=seconds. what is the maximum height?


1
Expert's answer
2021-02-24T06:54:29-0500

A function is a parabola with branches downward. A vertex has a co-ordinate

t=-80/(2 (-16))=2.5 then maximum height H=f(2.5)=-16 (2.5)2+80 (2.5)=100